suppose that 4000 is invested at an interest rate of 5.5% per year, compounded continuouly. what is the balance after 8 years?
Found 2 solutions by solver91311, stanbon:
Answer by solver91311(24713)


Use





The rest is calculator work.

John


Answer by stanbon(75887)

suppose that 4000 is invested at an interest rate of 5.5% per year, compounded continuouly. what is the balance after 8 years?
----------------
A(t) = P*e^(rt)
A(8) = 4000*e^(0.055*8) = $6210.83
=======================================
Cheers,
Stan H.
==============